Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
bersanetti, casanueva, derossi, gherardini, magazzu, mantovani, nardecchia, pinto, ruggi - 22:32 Wednesday 20 September 2023 (61729) Print this report
DAS differential tuning- CO2

Today, the  DAS differential tuning has been performed in order to reduce the current large value of BS offset needed to optimize the CMRF.

The procedure that we followed is listed below:

  1. lock at LOW NOISE 2   [not in low noise 3 because, in this condition, the diagonalization process is switched OFF]

  2. BS TY SET loop OFF 

  3. Waiting for the convergence of MICH/SRCL loop (~0.5 hrs) 

  4. BS scan in TY to evaluate the BS TY offset needed to minimize the CMRF.

Sequence of actions:

07:43 UTC: Lock at CARM NULL 1F

08.39 UTC: NI etalon control set to 20.25 C. 

08:47 UTC- 09.08UTC: OMC SCAN to take the reference for the starting point characterization 

09.31 UTC: Lock at LN2

10.09-10.15 UTC: WI DAS change (+15 % WI IN and +15% WI OUT wrt the starting point). WI inner and WI Outer in a row.

Some signals during the DAS change are shown in fig.1.

12.00-12.08 UTC:  the NI DAS has been changed (-15 % NI IN and -15% NI OUT wrt the starting point) during the unlock. 

The applied power changes on both NI and WI DAS are summarized in the table below:

 

STEP1_S

 

12.08 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.078*(1+15/100)=0.090

1.615*(1+15/100)=1.857

on the pickoff

0.418

0.0127*(1+15/100)= 0.0146

0.263*(1+15/100)=0.302

N

on the ITF

0.091

0.219*(1-15/100)+0.08=

0.266 

1.379*(1-15/100)+1.8 =2.972

on the pickoff

0.561

0.036*(1-15/100)+0.013=

0.044

0.225*(1-15/100)+0.293=

0.484

Then, the ITF unlocked 3 times between CARM NULL 1F and LN2 to the ongoing thermal transient causing a big divergence in the B1p PC 56 MHz upper and lower sidebands.

The first ‘good’ lock to evaluate the working point quality after the thermal transient, started at 13.27 UTC.

The CMRF appeared to be improved.

14.03 UTC:  BS TY scan started 

BS TY offset needed to minimize CMRF  is: -9e-4 (wrt -2e-3 as starting point)

14.20  UTC:  the BS TY offset has been re-settled to 0. 

Looking at the transient shown in fig.1, it seems that a good WP has been experienced during the increase of the WI inner ring power but followed by a fast degradation after the increase of the WI outer ring. Supposing  that the WP degradation has been caused by the increase of the WI outer ring, at 14.25 UTC, its power has been decreased by 15 % (PICKOFF INNER+OUTER=0.344W→0.292W) .

14.44 UTC : NI Set point [NI_RH_SET=20.2,rampTime=1]' sent to LSC_Etalon_Acl
14.44-UTC : WI Set point [WI_RH_SET=20.32,rampTime=1]' sent to LSC_Etalon_Acl

14.50 UTC - 15.06 UTC: Manuel and Paolo tested a new filter and performed noise injections (at the end of noise injection the ITF unlocked)

After that, we were not able to lock again in LN2 due to unbalance of the B1p PC 56 MHz. Camilla and Diego put some offsets on MICH (up to 15) but the problem persisted.
Thus, the last action on the WI outer has been reversed (PICKOFF INNER+OUTER=0.292W→0.344W) at 18.11 UTC in CARM NULL 1F to understand the effect of this ring on the unbalance. 

Unfortunately, at 18.16 UTC, the ITF unlocked.

At 18.43 UTC, the ITF relocked at CARM NULL 1F and then, we procedeed to LN2 without problems (no MICH OFFSET to balance the sidebands).

Between 19.38 UTC and 19.52 Camilla performed the BS TY scan and she found that the BS TY to minimize the CMRF is equal to -6.5e-4. 

Then, the BS TY offset has been re-settled to 0 in the time interval 19.53 UTC-19.58 UTC. 

A trend of all signals during the shift are shown in fig.2.

We left the ITF in LN3 (20.21 UTC), with BS TY SET loop closed and DASes power values identified as STEP1_S (the same step performed in August named STEP 2). 

Images attached to this report
Comments to this report:
derossi - 22:36 Wednesday 20 September 2023 (61733) Print this report

I put back the DARM line and OS cali in the ITF_lock.py lines #5586 and #5580.

I also re-engaged the BS TY SET loop (#5610 uncommented and #5878 commented.

mwas - 15:02 Thursday 21 September 2023 (61739) Print this report

Figure 1 shows the OMC scan performed yesterday.

Figure 2 shows for comparison the last OMC scan I could find, done on Jul 24.

On figure 1 the carier order 2 mode is a bight higher (30mW vs 20mW in July). The order 3-6 modes look the same, modes 7 and 8 are now much brighter by about a factor 2.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×